# A General-Purpose Multimodal Foundation Model for Dermatology

### v1

@@
A General-Purpose Multimodal Foundation Model for Dermatology
Siyuan Yan1,2, Zhen Yu1, Clare Primiero3, Cristina Vico-Alonso4, Zhonghua Wang1, Litao Yang1,2, Philipp Tschandl5, Ming Hu1,2, Gin Tan9, Vincent Tang10, Aik Beng Ng10, David Powell9, Paul Bonnington11, Simon See10, Monika Janda6, Victoria Mar4,8, Harald Kittler5, H. Peter Soyer∗3,7, Zongyuan Ge∗1,2
{affiliations}
@@ A General-Purpose Multimodal Foundation Model for Dermatology
Frazer Institute, The University of Queensland, Dermatology Research Centre, Brisbane, Australia.
@@
Department of Dermatology, Medical University of Vienna, Vienna, Austria.
@@
Dermatology Department, Princess Alexandra Hospital, Brisbane, Australia.
School of Public Health and Preventive Medicine, Monash University, Melbourne, Australia.
@@
*Corresponding author: Zongyuan Ge (zongyuan.ge@monash.edu) Peter Soyer (p.soyer@uq.edu.au)
Diagnosing and treating skin diseases require advanced visual skills across multiple domains and the ability to synthesize information from various imaging modalities. Current deep learning models, while effective at specific tasks such as diagnosing skin cancer from dermoscopic images, fall short in addressing the complex, multimodal demands of clinical practice. Here, we introduce PanDerm, a multimodal dermatology foundation model pretrained through self-supervised learning on a dataset of over 2 million real-world images of skin diseases, sourced from 11 clinical institutions across 4 imaging modalities. We evaluated PanDerm on 28 diverse datasets covering a range of clinical tasks, including skin cancer screening, phenotype assessment and risk stratification, diagnosis of neoplastic and inflammatory skin diseases, skin lesion segmentation, change monitoring, and metastasis prediction and prognosis. PanDerm achieved state-of-the-art performance across all evaluated tasks, often outperforming existing models even when using only 5-10% of labeled data. PanDerm’s clinical utility was demonstrated through reader studies in real-world clinical settings across multiple imaging modalities. It outperformed clinicians by 10.2% in early-stage melanoma detection accuracy and enhanced clinicians’ multi-class skin cancer diagnostic accuracy by 11% in a collaborative human-AI setting. Additionally, PanDerm demonstrated robust performance across diverse demographic factors, including different body locations, age groups, genders, and skin tones. The strong results in benchmark evaluations and real-world clinical scenarios suggest that PanDerm could enhance the management of skin diseases and serve as a model for developing multimodal foundation models in other medical specialties, potentially accelerating the integration of AI support in healthcare.
@@ Introduction
There is a pressing need to fully harness the potential of artificial intelligence (AI) in diagnosing and managing skin diseases. Although deep learning has demonstrated remarkable performance, often matching or surpassing dermatologists, current AI models for dermatology remain limited to isolated tasks, such as diagnosing skin cancer from dermoscopic images [1]. These models struggle to integrate diverse data types and imaging modalities, reducing their utility in real-world clinical settings. Dermatology, like internal medicine, is inherently complex, demanding a comprehensive, patient-centered approach. Diagnosing and treating skin cancer, for example, involves a range of tasks, including total body skin examination, risk assessment at both patient and lesion levels [2, 3, 4, 5], differentiation of neoplastic from inflammatory diseases [6], multimodal image analysis [7, 8], pathology interpretation [9, 10], monitoring lesion changes [11, 12], and predicting outcomes [13, 14]. The absence of integrated AI solutions capable of supporting these diverse workflows currently hampers the practical impact of AI in dermatology. Foundation models, however, hold the potential to address this gap by enabling a more holistic approach [15, 16].
@@
Foundation models are large-scale neural networks pretrained on vast, diverse datasets using self-supervised learning techniques, often leveraging weakly labeled or unlabeled data [17, 18, 19]. Built on rich knowledge representations, these models have demonstrated impressive performance across medical fields such as ophthalmology [20], radiology [21], and pathology [22, 23, 24, 25]. By leveraging large and diverse training datasets, these models are highly versatile in interpreting various data modalities [26], outperforming previous deep learning models in downstream tasks. Their strong feature representations also enable data-efficient applications [27, 28], requiring fewer labeled samples, which is a crucial advantage in domains where labeled data are scarce.
@@
The performance of foundation models is inherently linked to the scale of their parameters and training data [29, 30, 31]. In general computer vision, foundation models are pretrained on massive datasets like ImageNet [32] or JFT-300M [33] and most existing dermatology AI models still rely on these models for downstream adaptation. Some efforts have focused on self-supervised learning specifically for dermatology using public datasets [34, 35] or web-sourced skin images [36]. However, these approaches are often limited by dataset size, diversity, or the lack of real patient data, which hinders the ability of models to generalize effectively across the diverse tasks and modalities encountered in clinical dermatology.
@@
Here, we introduce PanDerm, a general-purpose, multimodal dermatology foundation model pretrained on over 2 million images sourced from 11 institutions across multiple countries, covering 4 imaging modalities (Fig 1a-c). This dataset represents the largest and most diverse image collection in dermatology to date. In the pretraining stage, PanDerm employs a novel combination of masked latent modeling and CLIP [37] feature alignment for self-supervised learning (Fig 1e and Methods), demonstrating superior data scalability and training efficiency compared to existing self-supervised algorithms (Fig 1f). The model exhibits advanced capabilities in holistic patient analysis by effectively interpreting a wide range of dermatological imaging modalities, including total body photography (TBP) as well as clinical, dermoscopic, and dermatopathology images. This versatility supports patient analysis throughout clinical workflows focused on, but not limited to, skin cancer detection (Fig 1d).
@@
We systematically evaluate PanDerm across 28 curated datasets (Fig 1g), covering a diverse array of clinical tasks, including screening, risk stratification, phenotype assessment, naevus counting, longitudinal monitoring, lesion change detection, diagnosis of neoplastic and inflammatory skin diseases, skin lesion segmentation, as well as recurrence prediction and prognosis. PanDerm achieves state-of-the-art performance on all tasks, often using only 5-10% of the labeled training data typically required. Furthermore, we demonstrate PanDerm’s clinical impact through three real-world studies, including malignant lesion detection using total body photography, human-AI collaboration, and early melanoma detection. These results collectively underscore PanDerm’s potential to enhance dermatological practice and may extend to other medical specialties, potentially accelerating the development and clinical integration of AI across various healthcare domains.
@@ Results
Figure 1: Overview of this study. Caption on next page.
(Previous page.) Figure 1: Overview of this study. a-c. The pretraining dataset includes 2.1 million images from 11 clinical sources and institutions across 4 modalities, representing the largest and most diverse multimodal dermatology dataset to date. Dataset distribution is shown by modalities (a), data sources (b), and institutions (c). d. PanDerm can interpret various imaging modalities and perform a wide range of dermatology-related clinical tasks. It has been evaluated in a real-world scenario study (e.g., TBP-based melanoma screening) and two reader studies (human-AI collaboration and early melanoma detection). Dermatopathology images: microscopic images of skin biopsy specimens at various magnifications. Clinical images: wide-field or regional images capturing the lesion and surrounding skin taken with digital cameras. Dermoscopic images: close-up images taken with a dermoscope, showing detailed lesion structures. TBP tiles: lesion crops taken from macro TBP images. e. PanDerm features a decoupled architecture comprising a ViT-large [38] encoder, a regressor, and a CLIP-based teacher model. Its pretraining process employs two objectives: a latent representation reconstruction objective and a CLIP latent alignment objective. f. PanDerm’s performance across varying pretraining data sizes and epochs, measured by average AUROC on 8 benchmark datasets, with comparative analysis against other pretraining strategies. g. PanDerm surpasses existing pretrained models on 28 evaluation datasets, covering various clinical tasks across 4 modalities.
Ablation studies and comparison with other self-supervised learning strategies
@@
To evaluate PanDerm’s effectiveness, we conducted ablation studies to examine the impact of pretraining data size and training epochs on downstream performance (datasets described in Extended Data Table 28). PanDerm demonstrated strong scalability, with consistent AUROC improvements as pretraining data increased from 0.8 to 1.8 million images (Fig 1f left). Notably, PanDerm outperformed SwAVDerm [36], a recent dermatology self-supervised learning model, using only 0.8 million images compared to SwAVDerm’s 1.2 million (Fig 1f left). Furthermore, PanDerm demonstrated superior training efficiency compared to state-of-the-art self-supervised learning models, achieving better performance with only 200 training epochs, while MILAN [39] required 500 epochs, and both DINOv2 [40] and MAE [19] needed 800 epochs (Fig 1f right), all using the same pretraining dataset. This efficiency stems from PanDerm’s novel use of CLIP as a teacher model for semantic feature learning. PanDerm also surpassed vision-language models such as CLIP [37], MONET [41], and BiomedCLIP [42] in benchmark evaluations (Extended Data Table 28). Additionally, It demonstrated emergent capabilities in dermatology similar to those of DINOv2 in natural images, with linear probing performance comparable to full-parameter fine-tuning (Extended Data Table 29). Building on these promising initial results, we extended our evaluation to a broader range of dermatological tasks, primarily comparing PanDerm with three representative pretrained models in the following sections: SL-Imagenet[32] and DINOv2[40] (both widely-used foundation models pretrained on natural images with a ViT-large [38] backbone), and SwAVDerm[36] (a self-supervised model pretrained on a large skin image dataset from search engines).
@@
We systematically evaluated PanDerm diagnostic performance across 10 public datasets from 4 imaging modalities and 7 international sites (Fig 2a). These datasets covered multi-class classification of pigmented neoplastic lesions and binary melanoma diagnosis tasks. We primarily evaluated performance using weighted F1 score for multi-class datasets and area under the receiver operating characteristic curve (AUROC) for binary class datasets. PanDerm consistently outperformed all other models, achieving significant improvements on 9 out of 10 datasets, with average gains of 5.1%, 8.0%, 4.2%, and 0.9% on dermoscopic, clinical, TBP, and pathology datasets, respectively (Fig 2a). On benchmarks like HAM10000 [35] and PAD [43], PanDerm surpassed the next-best models by 4.7% (P $<$ 0.001) and 9.0% (P $<$ 0.001), respectively (Fig 2a; Extended Data Table 1 and Extended Data Fig 3).
@@
When evaluating label efficiency generalization, PanDerm consistently outperformed other models across all datasets (Fig 2b, Extended Data Table 15-20). PanDerm matched the next-best model’s performance using only 10% to 30% of labeled data for skin lesion diagnosis across various modality datasets (Fig 2b), demonstrating PanDerm’s significant advantage in limited data scenario. Additional results for other tasks are presented in Extended Data Fig 4. To further assess PanDerm’s generalization ability, we extended a previous out-of-distribution generalization experiment [44, 45] to evaluate model performance for melanoma diagnosis on both dermoscopic and close-up clinical images from 7 external datasets. These datasets were collected from multinational institutions, representing populations distinct from the training data. PanDerm demonstrated significant superiority over all pretrained models, achieving higher AUROC scores across all external datasets (Fig 2c). Notably, PanDerm showed a significant improvement even on clinical images not used during fine-tuning, with AUROC gains of 4.0%, 2.6%, and 2.1% on the three clinical datasets (All P $<$ 0.001).
Figure 2: PanDerm’s versatile capacity in diverse diagnosis tasks. Caption on next page.
(Previous page.) Figure 2. PanDerm’s versatile capacity in diverse diagnosis tasks. a. Performance of PanDerm versus other pretrained models on 10 pigmented skin lesion datasets across multiple centers and modalities. n: data size, c: class number. Metrics: AUROC for binary class datasets (c=2) and W_F1 (weighted F1) score for multi-class datasets (c$>$2). Dashed lines indicate the average performance of each model across different datasets. b. Comparison between PanDerm and other pretrained models in a label efficiency generalization setting on four representative datasets, demonstrating model performance at various percentages of training data. Vertical dash lines show the data quantity needed for PanDerm to match existing model performance. c. External validation for melanoma diagnosis across 7 datasets. d. Performance evaluation of general skin condition classification (up to 74 classes) using clinical images. e. Early melanoma detection results (reader study 1) comparing PanDerm to 12 clinicians (7 experienced dermatologists, 5 registrars). X-axis: 89 melanoma lesion IDs; Y-axis: lesion image sequence length. Points on the histogram represent the initial time points of correct melanoma diagnoses. Points below y=0 correspond to melanoma lesions undetected throughout the sequence. Error bars in a, c, d show 95% CIs; bar centers in a, c, d represent mean value; dots in b represent mean value. Estimates are computed using nonparametric bootstrapping with 1000 bootstrap replicates. P-values are calculated using a two-sided t-test.
@@
We further evaluated PanDerm’s ability to classify a broader range of skin conditions, encompassing up to 74 classes. These included not only pigmented skin lesions but also inflammatory, infectious, and rare diseases. The evaluation was conducted on three clinical image datasets: the public DermNet dataset [46] (23 classes), and two in-house datasets, MMT-09 (9 classes) and MMT-74 (74 classes). As shown in Fig 2d, PanDerm achieved weighted F1 improvements of 3.2%, 7.1%, and 8.2% on MMT-09, DermNet, and MMT-74, respectively, compared to the next-best models (all P $<$ 0.001). We observed PanDerm’s performance advantage increased as the number of skin conditions grew, indicating a strong generalization and discriminative capacity across diverse diseases. PanDerm also outperformed all other pretrained models on all metrics across the three datasets (all P $<$ 0.001; Extended Data Table 2). In the DermNet dataset, PanDerm exceeded the next-best model’s area under the precision-recall curve (AUPR) by 14.7%. Further details on the experimental setup, datasets, and metrics are provided in Methods.
Reader study 1: Early melanoma detection compared with clinicians
We conducted a reader study to compare PanDerm’s performance in early melanoma detection using sequential images. We used a dermoscopic image dataset from Alfred Hospital [47], featuring multiple follow-up images of the same lesions over time. The study evaluated two key aspects: overall diagnostic accuracy and early melanoma detection capability. PanDerm’s performance was compared to that of 12 human reviewers (7 experienced dermatologists and 5 registrars). In terms of overall accuracy, PanDerm outperformed the average human reviewer by 10.2% and surpassed the best-performing human by 3.6%. For early detection, we assessed the earliest time point of correct malignant diagnosis for each melanoma case. PanDerm demonstrated superior ability in this challenging task, correctly identifying 77.5% (69 out of 89) of melanoma lesions at the first imaging time point, compared to only 32.6% (29 correct diagnoses) for human reviewers. Individual dots in the histograms of Fig 2e represent the earliest correct diagnosis time points for both PanDerm and human reviewers, visualizing the comparative early detection performance. Further details on the reader study setup and datasets are provided in Methods.
@@
Monitoring suspicious melanocytic lesions over a three-month period is a widely accepted procedure for early melanoma detection, as changes often prompt excision to rule out melanoma, whilst stability can be reassuring [12]. We assessed PanDerm’s ability for short-term lesion change detection by framing it as a binary classification task: determining whether a pair of short-term lesion images shows change or remains stable. A Siamese network architecture [48], which is well-suited for comparing image pairs to detect subtle changes, was employed. The primary challenges in this task are the subtlety of changes and the variability in imaging conditions. To address these challenges, we developed a comprehensive image pre-processing and alignment pipeline that includes dark corner removal, skin inpainting, image registration, and segmentation (Fig 3c). We then compared PanDerm’s performance on two longitudinal dermoscopic datasets, SDDI1 and SDDI2 (Fig 3a, Fig 3b, and Methods). Ablation studies (Fig 3d left) showed that our pre-processing pipeline significantly improved PanDerm’s performance on lesion monitoring, increasing AUROC from 0.596 (95% CI 0.567-0.624) to 0.706 (95% CI 0.686-0.725) in SDDI1 (P $<$ 0.001) and from 0.683 (95% CI 0.517-0.894) to 0.767 (95% CI 0.649-0.886) in SDDI2 (P $<$ 0.001). Using the optimized pipeline for all models, PanDerm achieved AUROC improvements of 4.3% in SDDI1 (P $<$ 0.001) and 3.7% in SDDI2 over the next-best model (Fig 3d, middle). For lesions later diagnosed as malignant (Fig 3b, middle), PanDerm achieved an AUROC of 0.840 (95% CI 0.769-0.911), surpassing the next-best model by 15.0% (P $<$ 0.01) (Fig 3d, right). Further details on lesion change detection are provided in Methods and Extended Data Table 3-5.
Figure 3: Short-term lesion change detection, metastasis prediction, and prognosis results. Caption on next page.
@@
(Previous page.) Short-term lesion change detection (a-d) and metastasis prognosis (e-k) results. a. SDDI1 dataset statistics for melanocytic lesion change detection: ratio of changed lesions during follow-up, ratio of changed malignant lesions during follow-up, and follow-up time distribution. b. Ratio of changed lesions in the SDDI2 dataset. c. Longitudinal dermoscopic image-based lesion change detection using PanDerm. For comparing subtle changes in paired lesions during short-term follow-up (e.g., 3 months), images undergo dark corner detection and removal, skin inpainting, registration, and lesion segmentation. This allows models to focus on subtle differences between lesions at different time points. d. Ablation study on pre-processing: “Default” (direct input of sequential images), “w/Warp” (registration only), “w/Mask” (lesion segmentation only), and “w/Whole pipeline”(complete pre-processing as in c). For change detection in all lesions and malignant lesions, all models are evaluated using the same whole pre-processing pipeline for fair comparison. e. Performance of binary metastasis prediction (control vs. metastasis), measured by AUROC. f. Scheme of PanDerm for predicting melanoma metastasis and prognosis. g. Distribution of metastasis types in the Combinemel dataset (MS represents metastasis). h. Kaplan–Meier curves for the recurrence-free interval (RFI) in invasive melanoma patients (CombinMel dataset), stratified by PanDerm prediction scores. i. Forest plots of hazard ratios for PanDerm, stratified groups in invasive melanoma patients. j. Time-dependent AUC of PanDerm vs. clinical variable score combinations. k. Time-dependent AUC comparison of PanDerm and other pretrained models. Error bars in d-e and j-k represent 95% CIs; bar centers indicate the mean value. Estimates are computed with five-fold cross-validation.

### v2

# A Multimodal Vision Foundation Model for Clinical Dermatology

@@
A Multimodal Vision Foundation Model for Clinical Dermatology
Siyuan Yan1,2, Zhen Yu1, Clare Primiero3, Cristina Vico-Alonso5, Zhonghua Wang1, Litao Yang1,2, Philipp Tschandl4, Ming Hu1,2, Lie Ju1, Gin Tan9, Vincent Tang10, Aik Beng Ng10, David Powell9, Paul Bonnington11, Simon See10, Elisabetta Magnaterra12, Peter Ferguson13,14, Jennifer Nguyen5, Pascale Guitera15, Jose Banuls16, Monika Janda6, Victoria Mar5,8†, Harald Kittler4†, H. Peter Soyer3,7†, Zongyuan Ge1†∗ {affiliations}
@@ A Multimodal Vision Foundation Model for Clinical Dermatology
Frazer Institute, The University of Queensland, Dermatology Research Centre, Brisbane, Australia
Department of Dermatology, Medical University of Vienna, Vienna, Austria
@@
Dermatology Department, Princess Alexandra Hospital, Brisbane, Australia
School of Public Health and Preventive Medicine, Monash University, Melbourne, Australia
@@
Section of Dermatology, Department of Health Sciences, University of Florence, Florence, Italy
Melanoma Institute Australia, North Sydney, New South Wales, Australia
Tissue Pathology and Diagnostic Oncology, Royal Prince Alfred Hospital and NSW Health Pathology, Sydney, New South Wales, Australia
Sydney Melanoma Diagnostic Centre, Royal Prince Alfred Hospital, New South Wales, Australia
Department of Dermatology, Hospital General Universitario de Alicante, ISABIAL, Alicante, Spain
[…]
Corresponding author: Zongyuan Ge (zongyuan.ge@monash.edu)
Diagnosing and treating skin diseases require advanced visual skills across domains and the ability to synthesize information from multiple imaging modalities. While current deep learning models excel at specific tasks like skin cancer diagnosis from dermoscopic images, they struggle to meet the complex, multimodal requirements of clinical practice. Here, we introduce PanDerm, a multimodal dermatology foundation model pretrained through self-supervised learning on over 2 million real-world skin disease images from 11 clinical institutions across 4 imaging modalities. We evaluated PanDerm on 28 diverse benchmarks, including skin cancer screening, risk stratification, differential diagnosis of common and rare skin conditions, lesion segmentation, longitudinal monitoring, and metastasis prediction and prognosis. PanDerm achieved state-of-the-art performance across all evaluated tasks, often outperforming existing models when using only 10% of labeled data. We conducted three reader studies to assess PanDerm’s potential clinical utility. PanDerm outperformed clinicians by 10.2% in early-stage melanoma detection through longitudinal analysis, improved clinicians’ skin cancer diagnostic accuracy by 11% on dermoscopy images, and enhanced non-dermatologist healthcare providers’ differential diagnosis by 16.5% across 128 skin conditions on clinical photographs. These results demonstrate PanDerm’s potential to improve patient care across diverse clinical scenarios and serve as a model for developing multimodal foundation models in other medical specialties, potentially accelerating the integration of AI support in healthcare.
@@ Introduction
There is a pressing need to fully harness the potential of artificial intelligence (AI) in diagnosing and managing skin diseases. Although deep learning has demonstrated remarkable performance, often matching or surpassing dermatologists, current AI models for dermatology remain limited to isolated tasks, such as diagnosing skin cancer from dermoscopic images [1]. These models struggle to integrate various data types and imaging modalities, reducing their utility in different real-world clinical settings. Dermatology, like internal medicine, is inherently complex, encompassing a broad spectrum of conditions from common dermatoses to life-threatening malignancies, necessitating a comprehensive, patient-centered approach that integrates various clinical workflows.
In clinical practice, diagnosing and treating skin conditions involves a range of tasks, including total body skin cancer detection and risk assessment [2, 3, 4, 5], differential diagnosis of hundreds of dermatological conditions such as inflammatory dermatoses and pigmentary disorders [6], multimodal image analysis [7, 8], pathology interpretation [9, 10], monitoring lesion changes [11, 12], and predicting outcomes [13, 14]. The absence of integrated AI solutions capable of supporting these various workflows currently hampers the practical impact of AI in dermatology. Recent advances in foundation models have emerged as a promising direction to address this challenge [15, 16].
@@
Foundation models are large-scale neural networks pretrained on vast, diverse data using self-supervised learning techniques, often leveraging weakly labeled or unlabeled data [17, 18, 19]. Built on rich knowledge representations, these models have demonstrated impressive performance across medical fields such as ophthalmology [20], radiology [21], and pathology [22, 23, 24, 25]. Through comprehensive pretraining on large and diverse data, these models develop versatile representations that can effectively adapt to various clinical scenarios, outperforming previous deep learning models in downstream tasks. Their strong feature representations also enable data-efficient applications [26, 27], requiring fewer labeled samples, which is particularly crucial for medical domains where expert-annotated data are often limited.
@@
However, developing effective foundation models for dermatology presents unique challenges. The performance of foundation models is inherently linked to the scale of their parameters and training data [28, 29, 30]. In general computer vision, foundation models are pretrained on massive datasets like ImageNet [31] or JFT-300M [32] and most existing dermatology AI models still rely on these models for downstream adaptation. Some efforts have focused on self-supervised learning specifically for dermatology using public datasets [33, 34] or web-sourced skin images [35]. However, these approaches are often limited by dataset size, diversity, or the lack of real patient data. Moreover, while recent advances in medical foundation models have shown promise in various specialties, they cannot fully address dermatology’s unique needs. Specialty-specific foundation models [23, 21, 20] typically focus on single imaging modalities, while general biomedical models, despite their broad scope, struggle with domain-specific data scarcity and integrating heterogeneous modalities for comprehensive clinical analysis.
Figure 1: Overview of this study. Caption on next page.
(Previous page.) Figure 1: Overview of this study. a-c. The pretraining dataset includes 2.1 million images from 11 clinical sources and institutions across 4 modalities. Dataset distribution is shown by modalities (a), data sources (b), and institutions (c). d. PanDerm can interpret various imaging modalities and perform a wide range of dermatology-related clinical tasks. It has been evaluated in a real-world scenario study (TBP-based melanoma screening), and three reader studies. Dermatopathology images: microscopic images of skin biopsy specimens. Clinical images: wide-field images capturing the lesion and surrounding skin. Dermoscopic images: close-up images taken with a dermoscope, showing detailed lesion structures. TBP tiles: lesion crops from macro TBP images. e. PanDerm features a decoupled architecture comprising a ViT-large [36] encoder, regressor, and CLIP-based teacher model, with pretraining using representation reconstruction and CLIP latent alignment objectives. f. PanDerm’s performance across varying pretraining data sizes and epochs, measured by average AUROC on 8 benchmark datasets, with comparative analysis against other strategies. g. PanDerm surpasses existing pretrained models on 28 evaluation datasets across 4 modalities.
@@
Here, we introduce PanDerm, a general-purpose, multimodal dermatology foundation model. Uniquely designed to integrate multiple imaging modalities, PanDerm is pretrained on over 2 million images sourced from 11 institutions across multiple countries, covering 4 imaging modalities spanning diverse dermatological conditions (Fig 1a-c). In the pretraining stage, PanDerm employs the masked latent modeling and CLIP [37] feature alignment for self-supervised learning (Fig 1e and Methods), demonstrating superior data scalability and training efficiency compared to existing self-supervised algorithms (Fig 1f). The model achieves unified representation learning across total body photography (TBP), clinical, dermoscopic, and dermatopathology images, enabling comprehensive patient analysis throughout diverse clinical workflows (Fig 1d).
@@
We systematically evaluate PanDerm across 28 benchmarks (Fig 1g), covering a diverse array of clinical tasks, including screening, risk stratification, phenotype assessment, naevus counting, longitudinal monitoring, lesion change detection, diagnosis of both common and rare skin conditions, skin lesion segmentation, as well as recurrence prediction and prognosis. PanDerm achieves state-of-the-art performance on all tasks, often using only 5-10% of the labeled training data typically required. Through three reader studies, we demonstrate that this unified multimodal approach outperforms clinicians in early-stage melanoma detection, enhances clinicians’ diagnostic accuracy in skin cancer diagnosis, and supports non-specialist healthcare providers in the differential diagnosis of diverse skin conditions. These findings highlight the potential of specialty-specific foundation models to advance medical practice by integrating diverse modalities, with broader implications for AI development across healthcare specialties.
@@ Results
Ablation studies and comparison with other training strategies
Figure 2: PanDerm’s versatile capacity in diverse diagnosis tasks. a. Performance comparison of PanDerm versus other pretrained models on 10 pigmented skin lesion datasets across multiple centers and modalities. n: data size, c: class number. Metrics: AUROC for binary class (c=2) and W_F1 score for multi-class (c$>$2) datasets. Dashed lines indicate average model performance across datasets. b. Comparison between PanDerm and other pretrained models in label efficiency generalization on four representative datasets, showing performance at various training data percentages. Vertical dash lines indicate the data quantity needed for PanDerm to match existing model performance. c. External validation for melanoma diagnosis across 7 datasets. d. Performance evaluation of general skin condition classification (up to 74 classes) using clinical images. Error bars in a, c, d show 95% CIs; bar centers in a, c, d represent mean value; dots in b represent mean value. Estimates were computed using nonparametric bootstrapping with 1000 bootstrap replicates. P-values calculated using a two-sided t-test.
@@
To evaluate PanDerm’s effectiveness, we conducted systematic analyses examining how model performance scales with training data and computational resources (datasets described in Extended Data Table 36). First, compared to existing dermatology-specific models, PanDerm showed strong scalability as training data increased from 0.8 to 1.8 million skin images (Fig 1f left). Notably, it achieved superior performance to SwAVDerm [35], a leading dermatology self-supervised learning model, using 33% less training data. When compared to other self-supervised training techniques, PanDerm demonstrated remarkable computational efficiency, requiring only 200 training epochs to achieve the best performance, compared to 500-800 epochs needed by leading methods like MILAN [38], DINOv2 [39] and MAE [19] (Fig 1f right). Furthermore, PanDerm also surpassed vision-language models such as CLIP [37], MONET [40], and BiomedCLIP [41] in benchmark evaluations (Extended Data Table 36), while showing emergent capabilities in dermatology similar to those of DINOv2 in natural images, with linear probing performance comparable to full-parameter fine-tuning (Extended Data Table 37). When evaluated against generalist biomedical foundation models, PanDerm demonstrated substantial advantages across different dermatological tasks. Compared to a representative model in this category, BiomedGPT [42], PanDerm showed 20.9% better AUROC in melanoma detection, 34.7% higher Weighted F1 in differentiating between skin conditions, and 19.6% improved weighted F1 in analyzing microscopic skin tissue images (Extended Data Table 38). Even using computationally efficient methods, PanDerm maintained its advantages, outperforming both linear-probe and fine-tune versions of BiomedGPT by 14.3% and 5.1% respectively in linear probing (Extended Data Table 39). Based on these promising results, we expanded our evaluation to compare PanDerm with three representative AI models: SL-Imagenet[31] and DINOv2[39] (both widely-used foundation models pretrained on natural images with a ViT-large [36] backbone), and SwAVDerm[35] (a self-supervised model pretrained on a large skin image dataset from search engines).
@@
We systematically evaluated PanDerm diagnostic performance across 10 public datasets from 4 imaging modalities and 7 international sites (Fig 2a). These datasets covered multi-class classification of pigmented neoplastic lesions and binary melanoma diagnosis tasks. PanDerm consistently outperformed all other models, achieving significant improvements on 9 out of 10 datasets, with average gains of 5.1%, 8.0%, 4.2%, and 0.9% on dermoscopic, clinical, TBP, and pathology datasets, respectively (Fig 2a). On representative dermoscopy and clinical benchmarks like HAM10000 [34] and PAD [43], PanDerm surpassed the next-best models by 4.7% (P $<$ 0.001) and 9.0% (P $<$ 0.001), respectively (Fig 2a; Extended Data Table 8 and Extended Data Fig 9).
@@
PanDerm demonstrated strong performance even with limited training data, achieving comparable results to other models while using only 10% to 30% of the labeled images (Fig 2b, Extended Data Table 22-27). Additional results for other tasks are presented in Extended Data Fig 10. To test PanDerm’s generalization applicability, we evaluated its performance on melanoma diagnosis using images from 7 external medical centers, representing patient populations different from the training data. PanDerm demonstrated significant superiority over all pretrained models, achieving higher AUROC scores across all external datasets (Fig 2c). Notably, it maintained high performance even on clinical photographs that were not used during training, with AUROC gains of 4.0%, 2.6%, and 2.1% on the three external clinical datasets (All P $<$ 0.001).
@@
Beyond skin cancer diagnosis, we evaluated PanDerm’s ability to diagnose a broader range of skin conditions commonly seen in clinical practice. We tested on three complementary datasets: the public DermNet dataset [44] covering 23 common conditions and two internal datasets - MMT-09 (9 conditions) and MMT-74 (74 conditions). These datasets comprehensively cover inflammatory diseases, infections, various types of skin tumors, and other frequently encountered skin problems. As shown in Fig 2d, PanDerm achieved weighted F1 improvements of 3.2%, 7.1%, and 8.2% on MMT-09, DermNet, and MMT-74, respectively, compared to the next-best models (all P $<$ 0.001). PanDerm’s advantage grew larger as the number of conditions increased, demonstrating its strong capability to handle complex, multi-disease scenarios. PanDerm also outperformed all other pretrained models on all metrics across the three datasets (all P $<$ 0.001; Extended Data Table 9). In the DermNet dataset, PanDerm exceeded the next-best model’s area under the precision-recall curve (AUPR) by 14.7%. Further details on the experimental setup, datasets, and metrics are provided in Methods.
@@
Monitoring suspicious melanocytic lesions over a three-month period is a widely accepted procedure for early melanoma detection, as changes often prompt excision to rule out melanoma, whilst stability can be reassuring [12]. We evaluated PanDerm’s ability to detect subtle changes in lesions over time by analyzing pairs of sequential dermoscopic images. To ensure accurate comparison despite variations in imaging conditions, we developed a comprehensive image processing system that standardizes image quality and alignment (Extended Data Fig 12). This processing system, combined with PanDerm’s advanced lesion change detection capabilities [45], significantly improved change detection accuracy from 0.596 (95% CI 0.567-0.624) to 0.706 (95% CI 0.686-0.725) in SDDI1 (Fig 3a; Fig 3c left) (P $<$ 0.001) and from 0.683 (95% CI 0.517-0.894) to 0.767 (95% CI 0.649-0.886) in SDDI2 (Fig 3b; Fig 3c left) (P $<$ 0.001). Using the optimized pipeline for all models, PanDerm achieved AUROC improvements of 4.3% in SDDI1 (P $<$ 0.001) and 3.7% in SDDI2 over the next-best model (Fig 3c, middle). For lesions later diagnosed as malignant, PanDerm achieved an AUROC of 0.840 (95% CI 0.769-0.911), surpassing the next-best model by 15.0% (P $<$ 0.01) (Fig 3c, right). Further details on lesion change detection method and dataset details are provided in Methods and Extended Data Table 10-12.
@@
Figure 3: Short-term lesion change detection and metastasis prognosis results. a. SDDI1 dataset statistics: ratio of changed lesions, the ratio of changed malignant lesions during follow-up, and follow-up time distribution. b. Ratio of changed lesions in SDDI2 dataset. c. Ablation study on pre-processing methods: Default” (direct input), w/Warp” (registration only), w/Mask” (lesion segmentation), and w/Whole pipeline” (complete pre-processing as in Extended Data Fig 12). For change detection, all models were evaluated using the whole pre-processing pipeline. d. Performance of binary metastasis prediction (control vs. metastasis) by AUROC. e. Scheme of PanDerm for melanoma metastasis and prognosis prediction. f. Distribution of metastasis types in Combinemel dataset (MS represents metastasis). g. Kaplan–Meier curves for the recurrence-free interval (RFI) in invasive melanoma patients (CombinMel dataset), stratified by PanDerm prediction scores. h. Forest plots of hazard ratios for PanDerm, stratified groups in invasive melanoma patients. i. Time-dependent AUC of PanDerm vs. clinical variable score combinations. j. Time-dependent AUC comparison of PanDerm and other pretrained models. Error bars in c-d and i-j represent 95% CIs; bar centers indicate the mean value. Estimates computed with five-fold cross-validation.